Three consecutive integers is 990. what is their sum? Show work.
Found 2 solutions by Alan3354, rashedbiman:
Answer by Alan3354(69443)

Three consecutive integers is 990.
990 is only one integer ???
----------------------
what is their sum?
It's "sum" is 990.
-------------------
Are you trying to trick us?

Answer by rashedbiman(2)

if the Three consecutive integers is 120. what is their sum?
Guys ths is very easy solution: good luck
ans.
120/2=60
again, 60/2=30
again, 30/2=15
again, 15/3=5
so here all the produnts is = 2*2*2*3*5
= 4*6*5
so the consecutive integers= 4*5*6
sum= 4+5+6
= 15
goodluck